# Planting and MCMC Sampling from the Potts model

### v1

@@ Appendix C Remaining Proofs for Section 4
We conclude (E.2) by noting $\mathbb{P}(X^{\prime}_{T_{1/4}}\neq\hat{X}_{kT_{1/4}}\mid{\mathcal{E}}_{<T^{\prime}})\leq\frac{\mathbb{P}(X^{\prime}_{T_{1/4}}\neq\hat{X}_{kT_{1/4}})}{\mathbb{P}({\mathcal{E}}_{<T^{\prime}})}=\frac{\mathbb{P}(X_{T_{1/4}}\neq\hat{X}_{T_{1/4}})}{\mathbb{P}({\mathcal{E}}_{<T^{\prime}})}$.
To finish, recall that $\mathbb{P}(\overline{{\mathcal{E}}_{<T^{\prime}}})\leq T^{\prime}\mathrm{e}^{-\Omega(n)}$ and $T_{1/4}=O(n\log n)$, thus there exists $\varepsilon_{0}=\mathrm{e}^{-\Omega(n)}$ such that $T_{1/4}\lceil\log_{2}(\varepsilon_{0}/2)\rceil\leq\tfrac{\varepsilon_{0}}{2}\leq\tfrac{1}{2}$. Then, by (E.3), for all $\varepsilon\in(0,\varepsilon_{0})$,

### v2

@@ Appendix E Remaining Proofs for Section 5
We conclude (F.2) by noting $\mathbb{P}(X^{\prime}_{T_{1/4}}\neq\hat{X}_{kT_{1/4}}\mid{\mathcal{E}}_{<T^{\prime}})\leq\frac{\mathbb{P}(X^{\prime}_{T_{1/4}}\neq\hat{X}_{kT_{1/4}})}{\mathbb{P}({\mathcal{E}}_{<T^{\prime}})}=\frac{\mathbb{P}(X_{T_{1/4}}\neq\hat{X}_{T_{1/4}})}{\mathbb{P}({\mathcal{E}}_{<T^{\prime}})}$.
To finish, recall that $\mathbb{P}(\overline{{\mathcal{E}}_{<T^{\prime}}})\leq T^{\prime}\mathrm{e}^{-\Omega(n)}$ and $T_{1/4}=O(n\log n)$, thus there exists $\varepsilon_{0}=\mathrm{e}^{-\Omega(n)}$ such that $T_{1/4}\lceil\log_{2}(\varepsilon_{0}/2)\rceil\leq\tfrac{\varepsilon_{0}}{2}\leq\tfrac{1}{2}$. Then, by (F.3), for all $\varepsilon\in(0,\varepsilon_{0})$,